let f(x) = (6x+3) and g(x)=(-2x+5)

*use chain rule to calculate h'(x) where h(x)=f(g(x)



1
Expert's answer
2021-09-20T11:04:51-0400

Given that"f(x)=(6x+3)" and "g(x)=(-2x+5)" , we can find "h'(x)" , where "h(x)=f(g(x))"


Now,

"h(x)=f(g(x)).........................................(1)"

Differentiate (1) with respect to x using chain rule we get

"h'(x)=\frac{d}{dx}[f(g(x))]"


"=f'(g(x)).g'(x)...........................................(2)"


Now f'(g(x) is the derivative of f with x replaced by g, this can be written as "\frac{df}{dg}"


Now,

"f(x)=(6x+3)"

so

"f'(x)=\frac{d}{dn}(6x+3)=6"

so

"f'(g(x))=6"

And

"g(x)=(-2x+5)"

so

"g'(x)=\frac{d}{dn}(-2x+5)"

"=-2"


so from (2)

"h'(x)=6\times-2=-12"